OpenStudy (anonymous):

384,400,000 in scientific notation

OpenStudy (anonymous):

3.844 x 10^8

OpenStudy (anonymous):

thanks, that's what I thought, it's just been so long since I've done this!

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

384,400,000 = 3844 * 10^5 = 3.844 * 10^8